OpenStudy (anonymous):

NEED HELP NOW!!! I WILL GIVE MEDALS!!!!!!!!!! What is the sum of x + y when x = –1.9 and y = –7.4?

OpenStudy (anonymous):

hallo do u like dogs

OpenStudy (anonymous):

hallo do u like dogs

OpenStudy (zzr0ck3r):

you are asking what \(-1.9 + (-7.4) = -1.9-7.4 = -(1.9+7.4) = -9.3\)

OpenStudy (nirmalnema):

-1.9+(-7.4)=-9.3

OpenStudy (anonymous):

No...I'm asking what is the sum of x + y when x = –1.9 and y = –7.4?

OpenStudy (zzr0ck3r):

correct the sum of x+y is x+y

OpenStudy (anonymous):

x+y=(-1.9)+(-7.4) =-1.9-7.4 =-9.3 hence the value of x+y is -9.3

OpenStudy (anonymous):

Wtf...

OpenStudy (anonymous):

HEY. do NOT cuss on OPENSTUFHY DOT COM.

OpenStudy (anonymous):

Huh? I didn't cuss..

OpenStudy (anonymous):

I would also like to know how to solve this problem...What is the sum of x + y when x = –1.9 and y = –7.4?

OpenStudy (zzr0ck3r):

It is sort of a weird way to ask a question. Summing something is the same thing as adding something. So you said how do I add x+y when x=... This is weird because you are saying the same thing twice, or you have left something out, like What is the sum of x+y and 13 when x = -1.9 and y = -7.4 then we would get -1.(-7.4)+ 13.

OpenStudy (zzr0ck3r):

that should say -1.9+(-7.4)+13

OpenStudy (anonymous):

Well what I did was I looked at it and replaced x with -1.9 and y -7.4, And I got a 100% so haha :P